OpenStudy (anonymous):

Toni rows a boat 4.5 km/h upstream and then turns around and rows 5.5 km/h back downstream to her starting point. If her total rowing time is 48 min, for how long does she row upstream? Express your answer to the nearest minute.

OpenStudy (phi):

 distance upstream equals distance downstream. Using tu for time upstream, and td for time downstream, and using rate x time= distance, you have 4.5*tu= 5.5*td also, the total time is tu+td= 48/60 (in hours! not minutes) replace td with td= 48/60 - tu 4.5*tu= 5.5*(4/5 - tu) where we simplified 48/60 to be 4/5 (divide top and bottom by 12) now solve for tu